# Purifying photon indistinguishability through quantum interference

### v1

@@ Abstract
In solid-state quantum emitters, photon distinguishability arises from either fast (compared to the photon lifetime) physical processes such as pure dephasing due to interactions with phonons in the environment (see Fig.1a), and slow processes (i.e. slower than the photon lifetime) such as spectral diffusion due to interactions with charges and polarization drifts [20, 11]. The standard approach to mitigating such noise contributions has so far been to reduce the temperature in order to reduce the number of phonons, develop ultra-low-noise chip devices with very low electrical noise, and to reduce the emission lifetime (via the Purcell effect) in order to reduce the interaction time with the acoustic phonon modes  [20]. Recent works have proposed a different mitigation strategy by using linear optical circuits and ancillary photons to purify single photons once they have been emitted  [21, 22]. Purification is a concept well-known in quantum information processing; it consists of consuming multiple copies of noisy quantum states to obtain a single output state where the noise is suppressed [23]. Its applications include the purification of entanglement [23, 24] and of “magic states” for universal fault-tolerant quantum computation [25]. Sparrow and Marshall have adapted this concept to purify single indistinguishable photons by processing multiple partially distinguishable and un-entangled photons through quantum interference in linear-optical circuits [21, 22].
@@
The success probability is 25% for the $n=2$ case and decreases exponentially for higher $n$. This decrease in success probability can be mitigated by changing the reflectivity of the final beam-splitter, and further optimized by using different interferometers with different heralding patterns [22]. Moreover, because the detection also provides a heralding of success, multiplexing techniques can then be used to turn the heralded probabilistic process into near-deterministic [27].
@@
Experimental setup. A schematic of the full experimental set-up, which implements two copies of the $n=2$ purification circuits and performs quantum interference between the two purified outputs, is shown in Fig. 2. Deterministic single-photon generation is performed with a QD embedded in photonic crystal waveguide (PCWs), detailed in the bottom left inset of Fig. 2. Electrical tuning of the QD, facilitated by low-noise electrical contacts, stabilizes the charge state, ensuring emission on the desired transition and minimizing spectral diffusion due to residual charge noise [28, 12]. The single photons are emitted in the PCW and fiber-coupled via a shallow-edged grating (SEG) and a cryo-compatible objective lens. An etalon can be employed as a frequency filter to optimize the indistinguishability of emitted photons by removing the undesired phonon-induced spectral sideband. Pumping the QD with resonant $\pi$-pulses at a repetition rate of 80 MHz yields single photons with a measured 16.1 MHz rate, corresponding to a 23.7% fiber-coupled efficiency of the single-photon source, and purity of $1-g^{(2)}(0)$ = (97.21 $\pm$ 0.01)%. The photon stream is then converted into four streams of simultaneous photons in different spatial modes through a time-to-space demultiplexing module. As shown in Fig. 2, it consists of three resonantly-enhanced electro-optic modulators (EOMs) and polarising beamsplitters (PBSs) arranged in a tree-like structure [29, 30] to route the photons in four different output modes and with different fiber lengths to compensate for the temporal delays. At this stage, we detect four-photon coincidences at a rate of 3.2 kHz. An additional free-space coupling is introduced before feeding into the fiber-based HOM interferometers to precisely control the polarization of photons via half- and quarter-wave plates (HWP and QWP, respectively).
@@
The purification circuits are implemented with optical fibers and involve two fiber beam-splitters (BSs) for each copy of the scheme, with the purified photons interfering in a final BS. Because the overall circuit is based on cascaded HOM interferometers, which are phase-insensitive, no active phase stabilization is required. The photons are finally directed to six superconducting-nanowire single-photon detectors (SNSPDs, 85% system efficiency) to measure the output configurations.
@@
Results. To analyse the improvement in indistinguishability induced by the purification, we first assess the raw HOM visibility by interfering only two demultiplexed photons (inputs 1 and 4 in Fig. 2) by blocking inputs 2 and 3. Inputs 2 and 3 are then unblocked to implement the full circuit to test quantum interference between purified photons and extract their purified indistinguishability (see Supplementary Information for details).
The purification protocol is tested in different experimental configurations of the QD, each with a different value of partial distinguishability between the emitted photons. The data from each configuration are shown in Fig. 3. In the first configuration (“No etalon” label in Fig. 3b), we test the protocol in a high-noise scenario by removing the etalon filter after the QD, which results in higher distinguishability due to the presence of incoherent spectral sidebands. This is indeed manifested in a measured low raw HOM visibility of $V=0.5829(1)$, as depicted in Fig. 3(a)(lower). When the purification protocol is implemented, the visibility of HOM interference between the purified photons is increased to $V=0.685(5)$ as depicted in Fig. 3(a) (upper), marking a significant improvement of $10.2(5)$%.
In a second QD configuration (“Optimal” label in Fig. 3b), the etalon is added to test the protocol in a low-noise environment where most phonon-induced noise is filtered out. The initial raw HOM visibility is measured to be $V_{0}=0.9050(1)$, which is then improved to $V_{f}=0.9327(1)$ using the purified photons. As discussed in the Supplementary Information, we estimate that in this case the remaining noise contributions limiting the visibility are dominated by spurious multi-photon terms arising by the finite $g^{(2)}(0)$ of the QD, while contributions due to partial distinguishability are mostly removed via the purification.
We test one additional scenario (“Sub optimal” label in Fig. 3b) where we keep the etalon but intentionally detune the voltage applied to the QD from its optimal value. Applying a voltage that is closer to the edge of the charge plateau increases the probability of co-tunelling of carriers to the contacts. This experimental condition induces a slight reduction in raw HOM visibility to $V_{0}=0.8332(1)$, which is then enhanced to a visibility of $V_{f}=0.9090(5)$ through purification.
@@
In Fig. 3 we also plot the theoretical estimations of the indistinguishability for the purified photons as a function of the raw visibilities, using three different models for predicting the partial distinguishability of the photons. The first model, which we call “multipermanent model” (dotted curve in Fig. 3b), assumes a constant state overlap and orthogonality between the distinguishable components of each photon, as outlined in [31]. A second model, based on the minimum purity model (MP, dashed curve in Fig. 3b) of Ref. [21], assumes that the internal degrees of freedom of all photons are in the same mixed state whose finite purity is the cause for distinguishability. Finally, we introduce a physically motivated model, denoted as the pure-dephasing model (PD, solid curve in Fig. 3b), that considers the noise processes inherent in a QD environment and incorporates higher-order indistinguishability instances (e.g., 3 and 4-photon contributions). The theoretical curves give slightly different expected improvements but are in good agreement with the observed measurements. For a comprehensive explanation and discussion of these diverse models, please refer to the Supplementary Information.
@@
After demonstrating successful purification for the case of a fast physical process, in particular phonon-induced dephasing, we test the purification protocol in the presence of slow errors as well. In this case, we deliberately altered the polarization of one input photon per purifier, as illustrated in Fig. 4a, and again evaluated the HOM visibility before and after purification. The results, shown in Fig. 4b, compare our experimental data (in purple) with a simulation illustrating the application of the purification protocol to photons with different polarizations. The polarization of the photons after going through multiple fiber BSs is transformed in randomized by temperature drift and bending of the fibers, and thus setting for a specific transformation is not attempted. The upper bound comes when the photons’ polarization is rotated in the same direction of the Bloch sphere, while the lower bound comes when they are rotated in opposite directions. Both cases are illustrated in the upper side of Fig. 4b, and the associated curves are plotted together with the measured raw and purified HOM visibilities. The observed data again show significant improvements in the indistinguishability and compatible with the theoretical bounds described above, demonstrating the realization of a successful purification protocol for both slow and fast noise contributions.
Conclusions. We have demonstrated a new type of purification process in quantum optical systems: the purification of partially distinguishable photons through quantum interference. Significant indistinguishability improvements are observed already for a small purification circuit involving $n=2$ input photons. We remark that the demonstrated capability to mitigate both fast and slow noise processes is in strong contrast to previous noise mitigation techniques, e.g. environment monitoring in solid-state emitters [32] and time-resolved measurement techniques [33, 34], which only work for slow noises (e.g. spectral diffusion and frequency mismatch in the above examples, respectively). The versatility of the approach highlights its relevance for any quantum photonic platform. In fact, the same approach can be straightforwardly used to purify photons also from other types of photon emitters, including atoms [7] and heralded photon sources [35].
@@
The purification protocols come at the cost of additional ancillary photons and, when relevant, multiplexing circuits to turn their probabilistic nature into near-deterministic. These are functionalities already required in photonic quantum computing architectures based on single photons [36], and indeed the hardware overhead of the purification protocols is likely largely dominated in practice by the overheads required in other parts of the architecture, such as in photonic entanglement generation circuits [37] and quantum error correction [38, 39]. Furthermore, the purification circuits implemented here can be significantly improved by modulating the reflectivities of the BSs or using discrete Fourier-transform interferometers [21, 22], enabling significantly higher success probabilities and indistinguishability improvements. Our study provides new and versatile experimental tools to purify photons and mitigate fundamental noise processes that ultimately may limit the scaling-up of photonic quantum technologies based on quantum emitters.
@@
Acknowledgements We thank Alex E. Jones for fruitful discussions. We acknowledge funding from the Danish National Research Foundation (Center of Excellence “Hy-Q,” grant number DNRF139), the Novo Nordisk Foundation (Challenge project ”Solid-Q”), the European Union’s Horizon 2020 research and innovation program under Grant Agreement No. 820445 (project name Quantum Internet Alliance) and No. 899368 (EPIQUS), the Marie Skłodowska-Curie grant agreement No 956071 (AppQInfo), and the QuantERA II Programme under Grant Agreement No 101017733 (PhoMemtor); from the Austrian Science Fund (FWF) through 10.55776/COE1 (Quantum Science Austria), 10.55776/F71 (BeyondC) and 10.55776/FG5 (Research Group 5); from the Austrian Federal Ministry for Digital and Economic Affairs, the National Foundation for Research, Technology and Development and the Christian Doppler Research Association. S.P. acknowledges financial support from the European Union’s Horizon 2020 Marie Skłodowska-Curie grant No. 101063763, from the Villum Fonden research grants No.VIL50326 and No.VIL60743, and from the NNF Quantum Computing Programme.

### v2

@@ Abstract
In solid-state quantum emitters, photon distinguishability arises from either fast (compared to the photon lifetime) physical processes such as pure dephasing due to interactions with phonons in the environment (see Fig.1a), and slow processes (i.e. slower than the photon lifetime) such as spectral diffusion due to interactions with charges and polarization drifts [20, 11]. The standard approach to mitigating such noise contributions has so far been to reduce the temperature in order to reduce the number of phonons, develop ultra-low-noise chip devices with very low electrical noise, and to reduce the emission lifetime (via the Purcell effect) in order to reduce the interaction time with the acoustic phonon modes  [20].
Recent works have proposed a different mitigation strategy by using linear optical circuits and ancillary photons to purify single photons once they have been emitted [21, 22]. Purification is a well-known concept in quantum information processing; it consists of consuming multiple copies of noisy quantum states to obtain a single output state where the noise is suppressed [23]. Its applications include the purification of entanglement [23, 24] and of “magic states” for universal fault-tolerant quantum computation [25]. Sparrow and Marshall have adapted this concept to purify single indistinguishable photons by processing multiple partially distinguishable and un-entangled photons through quantum interference in linear-optical circuits [21, 22].
@@
The success probability is 25% for the $n=2$ case and decreases exponentially for higher $n$.
This decrease can be mitigated by changing the reflectivity of the final beam-splitter, and further optimized by using different interferometers with different heralding patterns [22]. Moreover, because the detection also provides a heralding of success, multiplexing techniques can then be used to turn the heralded probabilistic process into near-deterministic [27].
@@
Experimental setup. A schematic of the full experimental set-up, which implements two copies of the $n=2$ purification circuits and performs quantum interference between the two purified outputs, is shown in Fig. 2 Deterministic single-photon generation is achieved from a neutral exciton of an InAs QD embedded in photonic crystal waveguide (PCW), detailed in the bottom left inset of Fig. 2. The QD is pumped resonantly with a pulsed laser, spectrally shaped with a home-built folded 4-f system, to set a bandwidth of $\sim$90 pm and match the QD wavelength $\lambda$=938.4 nm. Electrical tuning of the QD, facilitated by low-noise electrical contacts, stabilizes the charge state, ensuring emission on the desired transition and minimizing spectral diffusion due to residual charge noise [28, 12]. The single photons are emitted in the PCW and fiber-coupled via a shallow-edged grating and a cryo-compatible objective lens. An etalon with a bandwidth of 32 GHz can be employed as a frequency filter to optimize the indistinguishability of emitted photons by removing the undesired phonon-induced spectral sideband. Pumping the QD with resonant $\pi$-pulses at a repetition rate of 80 MHz yields single photons at a measured rate of 16.1 MHz (with an 85% efficient detection system), corresponding to a 23.7% fiber-coupled efficiency of the single-photon source, and purity of $1-g^{(2)}(0)$ = (97.21 $\pm$ 0.01)%. The photon stream is then converted into four streams of simultaneous photons in different spatial modes through a time-to-space demultiplexing module. As shown in Fig. 2, it consists of three resonantly-enhanced electro-optic modulators (EOMs) and polarising beamsplitters (PBSs) arranged in a tree-like structure [29, 30] to route the photons in four different output modes and with different fiber lengths to compensate for the temporal delays. At this stage, we detect four-photon coincidences at a rate of 3.2 kHz. An additional free-space coupling is introduced before feeding into the fiber-based HOM interferometers to precisely control the polarization of photons via half- and quarter-wave plates (HWP and QWP, respectively).
@@
The purification circuits are implemented with optical fibers and involves two fiber beam-splitters (BSs) for each copy of the scheme, with the purified photons interfering in a final BS. We characterized the optical losses induced by the purification setup to be of 1 dB. These are mainly due to fiber connectors between the beam splitters. Because the overall circuit is based on cascaded HOM interferometers, which are phase-insensitive, no active phase stabilization is required. The photons are finally directed to six superconducting-nanowire single-photon detectors (SNSPDs, average 85% system efficiency) to measure the output configurations.
@@
Results. To analyse the improvement in indistinguishability induced by the purification, we first assessed the raw HOM visibility by interfering only two demultiplexed photons (inputs 1 and 4 in Fig. 2) by blocking inputs 2 and 3. Inputs 2 and 3 are then unblocked to implement the full circuit to test quantum interference between purified photons and extract their purified indistinguishability (see Supplementary Information for details).
@@
The purification protocol was tested in different experimental configurations of the QD, each with a different value of partial distinguishability between the emitted photons. The data from each configuration are shown in Fig. 3. In the first configuration (“No etalon” label in Fig. 3b), we test the protocol in a high-noise scenario by removing the etalon filter after the QD, which results in higher distinguishability due to the presence of incoherent spectral sidebands. This is indeed manifested in a measured low raw HOM visibility of $V_{0}=0.5829(1)$, as depicted in Fig. 3(a)(lower). When the purification protocol is implemented, the visibility of HOM interference between the purified photons was increased to $V_{f}=0.685(5)$ as depicted in Fig. 3(a) (upper), marking a significant improvement of $10.2(5)$%. The purple curve is both narrower and has a lower maximum than the red curve, suggesting that both the dynamics and the amplitude of the distinguishable components are altered. The width of the peak nonetheless only changes due to the lower heralding probability of photons far away from the center position, as the effect can also be seen on the sidepeaks. It can thus not be concluded that the purified photons have different properties than the raw ones, except for a lower distinguishable component.
In a second QD configuration (“Optimal” label in Fig. 3b), the etalon was added to test the protocol in a low-noise environment where most phonon-induced noise is filtered out. The initial raw HOM visibility was measured to be $V_{0}=0.9050(1)$, which is then improved to $V_{f}=0.9327(1)$ using the purified photons. As discussed in the Supplementary Information, we estimate that in this case the remaining noise contributions limiting the visibility are dominated by spurious multi-photon terms arising by the finite $g^{(2)}(0)$ of the QD, while contributions due to partial distinguishability is mostly removed via the purification.
We tested one additional scenario (“Sub optimal” label in Fig. 3b) where we keep the etalon but intentionally detune the voltage applied to the QD from its optimal value. Applying a voltage that is closer to the edge of the charge plateau increases the probability of co-tunelling of carriers to the contacts. This experimental condition induced a slight reduction in raw HOM visibility to $V_{0}=0.8332(1)$ , which is then enhanced to visibility of $V_{f}=0.9090(5)$ through purification.
The data histograms which are the base for the Sub-optimal and Optimal configuration results are shown in the supplementary material. In Fig. 3 we also plot the theoretical estimations of the indistinguishability for the purified photons as a function of the raw visibilities, and used in addition three different models for predicting the partial distinguishability of the photons. The first model, which we call “multipermanent model” (dotted curve in Fig. 3b), assumes a constant state overlap and orthogonality between the distinguishable components of each photon, as outlined in [31]. A second considered model, based on the minimum purity model (MP, dashed curve in Fig. 3b) of Ref. [21], assumes that the internal degrees of freedom of all photons are in the same mixed state whose finite purity is the cause for distinguishability. Finally, we introduce a physically motivated model, denoted as the pure-dephasing model (PD, solid curve in Fig. 3b), that considers the noise processes inherent in a QD environment and incorporates higher-order indistinguishability instances (e.g., 3 and 4-photon contributions). The theoretical curves give slightly different expected improvements but are in good agreement with the observed measurements. For a comprehensive explanation and discussion of these diverse models, please refer to the Supplementary Information.
@@
After demonstrating successful purification for the case of a fast physical process, in particular phonon-induced dephasing, we test the purification protocol in the presence of slow errors as well. In this case, we deliberately altered the polarization of one input photon per purifier, as illustrated in Fig. 4a, and again evaluated the HOM visibility before and after purification. The results, shown in Fig. 4b, compare our experimental data (in purple) with a simulation illustrating the application of the purification protocol to photons with different polarizations. The polarization of the photons after going through multiple fiber BSs is transformed in randomized by temperature drift and bending of the fibers, and thus setting for a specific transformation is not attempted. The upper bound comes when the photons’ polarization is rotated in the same direction of the Bloch sphere, while the lower bound comes when they are rotated in opposite directions. Both cases are illustrated in the upper side of Fig. 4b, and the associated curves are plotted together with the measured raw and purified HOM visibilities. Representative histograms used to generate three of the data-points from the figure are shown in the supplementary material. The observed data again show significant improvements in the indistinguishability and compatible with the theoretical bounds described above, demonstrating the realization of a successful purification protocol for both slow and fast noise contributions.
It is important to mention that when comparing success probabilities of individual instances, spectral filtering can give a better performance than purification. However, there are essential differences between the two approaches that result in different advantages. First, filtering is a loss process. Its probabilistic nature is fundamental and there is no way to recover it as a near-deterministic operation. In contrast, our approach can be implemented through multiplexing of heralded events. Secondly, spectral filtering is effective in reducing distinguishability for certain noise sources (e.g. spectral detuning, phonon sidebands, spectral correlations in spontaneous photon sources) but is ineffective in correcting most incoherent processes, such as pure dephasing. Remarkably our scheme is not facing this limitation, making it operationally distinct from spectral filtering by being able to purify also incoherent processes.
@@
Conclusions. We have demonstrated a new type of purification process in quantum optical systems: the purification of partially distinguishable photons through quantum interference. Significant indistinguishability improvements are observed already for a small purification circuit involving $n=2$ input photons. We remark that the demonstrated capability to mitigate both fast and slow noise processes is in strong contrast to previous noise mitigation techniques, e.g. environment monitoring in solid-state emitters [32] and time-resolved measurement techniques [33, 34], which only work for slow noises (e.g. spectral diffusion and frequency mismatch in the above examples, respectively). The versatility of the approach highlights its relevance for any quantum photonic platform. In fact, the same approach can be straightforwardly used to purify photons also from other types of photon emitters, including atoms [7] and heralded photon sources [35]. The purification protocols come at the cost of additional ancillary photons and, when relevant, multiplexing circuits to turn their probabilistic nature into near-deterministic. These are functionalities already required in photonic quantum computing architectures based on single photons [36]. The hardware overhead of the purification protocols is anyway likely largely dominated in practice by the overheads required in other parts of the architecture, such as in photonic entanglement generation circuits [37] and quantum error correction [38, 39]. Furthermore, the purification circuits implemented here can be significantly improved by modulating the reflectivities of the BSs or using discrete Fourier-transform interferometers [21, 22], enabling significantly higher success probabilities and indistinguishability improvements. Our study provides new and versatile experimental tools to purify photons and mitigate fundamental noise processes that ultimately may limit the scaling-up of photonic quantum technologies based on quantum emitters.
@@
Acknowledgements We thank Alex E. Jones for fruitful discussions. We acknowledge funding from the Danish National Research Foundation (Center of Excellence “Hy-Q,” grant number DNRF139), the Novo Nordisk Foundation (Challenge project ”Solid-Q”), the European Union’s Horizon 2020 research and innovation program under Grant Agreement No. 820445 (project name Quantum Internet Alliance) and No. 899368 (EPIQUS), the Marie Skłodowska-Curie grant agreement No 956071 (AppQInfo), and the European Union’s Horizon Europe research and innovation program under Grant Agreement No. 101135288(EPIQUE) and No. 101017733 (QuantERA II Programme, project PhoMemtor). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council executive Agency. This research was funded in whole or in part by the Austrian Science Fund (FWF) through 10.55776/COE1 (Quantum Science Austria), 10.55776/F71 (BeyondC) and 10.55776/FG5 (Research Group 5); from the Austrian Federal Ministry for Digital and Economic Affairs, the National Foundation for Research, Technology and Development and the Christian Doppler Research Association. S.P. acknowledges financial support from the European Union’s Horizon 2020 Marie Skłodowska-Curie grant No. 101063763, from the Villum Fonden research grants No.VIL50326 and No.VIL60743, and from the NNF Quantum Computing Programme.
@@ Appendix C Data analysis
Appendix D Supplementary data
The histograms of data used to generate figure 3b in the main text can be seen in figure A4, where the no-etalon histogram is the one already presented. We also provide some representative histograms for the data used to generate figure 4b in the main text in figure A5
Figure A4: Histograms used to generate the figure 3b in the main text.
Figure A5: Representative histograms used to generate the figure 4b in the main text. (3rd, 5th and 9th datapoints)